OpenStudy (anonymous):

rationalize sqr(2)/3sqr(3)-sqr(2)

OpenStudy (anonymous):



OpenStudy (campbell_st):

multiply by 1 in the form \[\frac{3\sqrt{3}+\sqrt{2}}{3\sqrt{3}+\sqrt{2}}\] this creates the difference on 2 squares in the denominator and will leave a rational number.. you need to evaluate \[\frac{\sqrt{2}}{3\sqrt{3}-\sqrt{2}} \times \frac{3\sqrt{3}+\sqrt{2}}{3\sqrt{3}+\sqrt{2}}\]

OpenStudy (anonymous):



OpenStudy (campbell_st):

\[3\sqrt{3} \times 3\sqrt{3} = (3\times3)(\sqrt{3}\times \sqrt{3}) = 9 \times3 = 27\] \[\sqrt{4} = 2\] so you denominator should be 27 - 2 = 25